# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 3.3 Robustness with Comparable Accuracy
We first provide a preliminary understanding about the robustness of FL model. We train centralized model for 200 epochs and federated model for 400 rounds resulting in a decent accuracy of over 90% (see the regular column of Table 1). For CNN model, we train 200 epochs for centralized model and 600 rounds for federated model to achieve and accuracy of over 75%.
Table 1: Centralized and federated model under white-box attack.
@@
We can observe that both clean and adversarial accuracy of federated model is lower than its centralized counterpart, aligned as the result in zizzo2020fat . However, we conjecture that such an increase in adversarial accuracy is not attributed to the intrinsic robustness of centralized model but largely due to its high clean accuracy. To validate this hypothesis and facilitate a fair comparison between the two paradigms, we early-stop both models when their clean accuracy reach 90% (75% for CNN) and report the results in the same-acc column of Table 1. We early stop at 80% for adversarial training (72% for CNN). We can see, when both models reach a comparable clean accuracy, federated model shows greater robustness against white-box attack compared with the centralized model.
4 Experiments
After having some preliminary understandings regarding the robustness of FL under white-box attack, we discuss the transfer attack with limited data in this section. As described in Section 1, a much more realistic setting is that a group of hostile clients lay benign during training and leverage the fraction of training data to perform transfer attack. Consequently, the robustness against transfer attack of the FL model becomes vital in evaluating the security and robustness of FL applications. In this section, we first discuss the transfer attack in Section 4.1 and then explore the possibility of attacking with limited data in Section 4.2.
@@ 4.1 Robustness Against Transfer Attack
We explore the examples generated by two different training paradigms and their transferability when applying them to different models. Since similarity of decision boundary and clean accuracy influences and reflects the transferability between models goodfellow2014explaining ; liu2016delving ; demontis2019adversarial , we early stop both federated and centralized models at 90% of accuracy (75% for CNN). Note that we follow this training setting for the rest of this paper.
@@
From Table 2, we can see that the adversarial examples generated by federated model are highly transferable to both federated and centralized model while adversarial examples generated by centralized model exhibit less transferability. The T.Rate of federated-to-centralized attack is even larger than centralized-to-centralized attack. Secondly, T.Rate of adversarial examples between models trained under same paradigms is larger than models trained under different paradigms, which can be attributed to the difference of the two training paradigms, e.g. the discrepancy in the decision boundary goodfellow2014explaining ; liu2016delving or different sub-space tramer2017space .
@@ 4.2 Transfer Attack with limited data
In this section, we will evaluate the possibility of our proposed attack setting, i.e. transfer attack with limited data. To simulate this scenario, we fixed the generated partition used in the federated training and randomly selects a specified number of users as malicious clients whose data is available for performing the attack. To perform the transfer attack, we trains a substitute model in an centralized manner with the collected data. Training details are specified in Section 3.1. One of the key differences of proposed setting and conventional transfer-based attack is number of available data to train the substitute (source) model. The number of clients used to train the source model also serves as a key factor for a successful attack since one would reason that more data will lead to higher success rate. To provide an overview of the transferability of adversarial examples generated by the source model trained with different number of clients, we plot the their relation in Figure 1. We have the following observations:
@@
From observation 2, we can see that the proposed attack can achieve comparable or even better T.Rate or lower T.Acc. Consequently, we can conclude that the proposed attack setting with limited data is likely to cause significant security breaches in the current and future FL applications. To further explain observation 1 and an intriguing phenomenon that the T.Rate of ResNet50 model rises to the peak and then decreases, we hypothesize that when the number of clients used to train the source model is small, the clean accuracy of the source model is also low, leading to a large discrepancy in the decision boundary. Increasing the number of users used in the source model minimizes such discrepancy until the amount of data is sufficient to train a source model with similar accuracy. At this point, the difference between the federated and centralized caldarola2022improving becomes the dominant factor affecting the transferability since the source model is trained in centralized paradigm.
@@
Training substitute model in federated manner. To testify the conjecture mentioned above, we train the substitute model with the same number of clients’ data in federated manner. Specifically, since we have the knowledge of each samples’ clients, we partition the collected data from malicious clients as the federated model trains the source model in federated manner. We discuss the importance of such knowledge to transfer attack in Appendix E. Results can be seen in (c) of Figure 1. We can see that, with federated source model, the T.Rate can be slightly boosted at the beginning (limited number of clients) but continue to increase as the number of users increases and finally contributes to a significantly high T.Rate of 99%. This demonstrates our conjecture regarding the trend and also shows that if the hostile party trains the substitute model in federated manner, the T.Rate can be further increased despite of its computation burden.
@@
Boosting transfer rate. Following this conjecture, we first boost the T.Rate source model trained with limited data (e.g. 5 or 7 users) with data augmentation and model pretraining. Without loss of generality, in this paper, we leverage AutoAugment cubuk2018autoaugment and ImageNet pretrain. From (c) of Figure 1, we can clearly see, with these techniques we can successfully increase the T.Rate of 1% and 2% of clients from around 3% to 36% and 48% respectively. With 7% or 10% of clients’ data, the proposed attack setting achieves a high T.Rate of more than 80% (10% higher than the transfer-based attack). That is to say, with simple training techniques, malicious clients can attack with more than 40% of success rate with one or clients and 80% with 7 to 10 clients. To further illustrate the threat posed by our attack, we leverage the more advanced attack to replace the default PGD attack as shown in (c) of Figure 1. We can see that both auto-attack croce2020reliable and skip attack wu2020skip achieve a transfer rate of over 80% with only 3 and 10 users respectively. Noticeably, auto-attack achieves the highest transfer rate of 85% with 10% of the users. These results further demonstrate the threat our attack poses.
@@
Practical evaluation. To practically evaluate our attack setting, we consider two cases: 1. the malicious clients can only participate in some periods and 2. the attacker has no knowledge of the model deployed. We simulate these situations by attacking in different training stages and with different architectures, as shown in (d) of Figure 1 where a similar trend mentioned above is observed. We also conduct experiments with 10 and 30 users as shown in (e) of Figure 1 which we believe is a relatively simple setting that can further emphasize the threat of our attack setting.
Figure 1: Attack with data from a limited number of users.
@@ 5 Two intrinsic properties contributing to transfer robustness
To fully understand how adversarial examples transfer between centralized and federated models, we study two intrinsic properties of FL and its relation with transfer robustness. To protect the privacy of clients and leverage the massive data from user-end, FL utilizes distributed data to train a global model through local updates and aggregation at the server mcmahan2017communication . As a consequence, the distributed data and the aggregation operation is the core component of a FL method.
Figure 2: T.Rate vs. data of different heterogeneity and dispersion degree. (a): the top 3 are results of ResNet50; (b): the bottom 3 are results of CNN; Left: T.Rate as a function of the number of users in federated training; Middle: T.Rate as a function of dirichlet alpha; Right:T.Rate as a function of unbalanced sgm.
@@ 5.1 Decentralized training and data Heterogeneity
This section aims to explore the relationship between the degree of degree and transfer robustness.
@@
Control the degree of dispersion and heterogeneity. To explore the impact of distributed data on adversarial transferability, we control decentralization and heterogeneity through four indexes. By varying the number of clients in the partition, we alter the degree of dispersion of distributed data. For heterogeneity, we change the number of maximum classes per client mcmahan2017communication , the alpha value of Dirichlet distributions ($\alpha$) wang2020federated ; yurochkin2019bayesian (smaller $\alpha$ means a more non-iid partition) and the log-normal variance (sgm) of the Log-Normal distribution (larger variance denotes more unbalanced partitions) used in unbalanced partition zeng2021fedlab . We leverage FedLab framework to generate the different partitions zeng2021fedlab . For the rest experiments, we leverage the centralized trained model as the source to perform the transfer attack.
Degree of decentralization reduces transferability. As seen in Figure 2 (left of (a) and (b)), we can observe that T.Rate drops with an increasing number of users, which demonstrate that more decentralized data leads to lower transferability.
Data heterogeneity affects transfer attack. As per the middle plot of (a) and (b) in Figure 2, we can see that T.Rate increases clearly as $\alpha$ increases. We report the results on different the number of maximum classes per client in Appendix B. We also explore whether unbalanced data leads to a decrease of transferability in Figure 2 (right of (a) and (b)), where larger variance leads to a lower transfer rate. To validate the above observation, we provide statistical testing for the correlation coefficient in Appendix C. With the Spearman correlation coefficient, we report a significant negative correlation on ResNet50 between log-normal variance and T.Rate under a significance level of 0.1 ($p$-value=0.05). We report a significant correlation on all results with a level of 0.05 except the CNN experiments with different Dirichlet $\alpha$ and unbalanced sgm. We also visualize the linear regression results in Appendix I. Practically, we also evaluate this relation with more realistic attack settings. By attacking in different training stages and with different architectures as shown in (c) and (d) of Figure 2, we demonstrate that this correlation holds in various configurations. To sum up, we demonstrate that decentralized training reduces the transferability of adversarial examples. This can be attributed to the enlarged discrepancy of decision boundary between the source model and target model when the training data is more decentralized or heterogeneous. The heterogeneous distribution leads to a loss landscape significantly different from the ones of iid and centralized model and also results in a loss landscape with high variance caldarola2022improving , which all contribute to lower the transfer possibility of adversarial examples demontis2019adversarial .

### v2

@@ 4.1 Robustness with Comparable Accuracy
Table 1: Centralized and federated model under white-box attack. We can see that, with comparable clean accuracy, the FL model shows greater robustness against white-box attacks compared with the centralized model. This observation is consistent across different datasets and model architectures.
@@
ImageNet200
centralized R50 55.05 3.68 65.79 8.41
R50 (adv) 50.04 31.20 55.59 32.84
federated R50 55.03 13.42 60.59 15.68
R50 (adv) 50.18 38.31 54.92 41.13
In order to provide a preliminary understanding about the robustness of the FL model, we train the centralized model for 200 epochs and the federated model for 400 rounds resulting in a decent accuracy of over 90% (see the regular column of Tab. 1). For the CNN model, we train 200 epochs for the centralized model and 600 rounds for the federated model to achieve an accuracy of over 75%.
We can observe that the federated model’s clean and adversarial accuracy is lower than its centralized counterpart, aligned as the result in (Zizzo et al., 2020). However, we conjecture that such an increase in adversarial accuracy is not attributed to the intrinsic robustness of the centralized model but largely due to its high clean accuracy. To validate this hypothesis and facilitate a fair comparison between the two paradigms, we early-stop both models when their clean accuracy reaches 90% (75% for CNN) and report the results in the same-acc column of Tab. 1. We early stop at 80% for adversarial training (72% for CNN). We can see that when both models reach a comparable clean accuracy, the FL model shows greater robustness against white-box attacks compared with the centralized model.
To further validate our hypothesis, we perform the experiment on a much larger and more realistic dataset, i.e. ImageNet200. We early stopped both models at 55% accuracy. Results can be seen in the bottom two rows of Tab. 1. We can see that FL models demonstrate superior robustness against white-box attacks compared with the centralized model on both same-acc and regular settings.
@@ 4.2 Robustness Against Transfer Attack
ImageNet200
R50 federated 2.29 / 95.60 6.52 / 86.74
centralized 22.72 / 54.00 8.27 / 83.13
We turn to the black-box attack which is more practical and realistic in real-world applications. We explore the examples generated by two different training paradigms and their transferability to different models. Since the similarity of decision boundary and clean accuracy influences and reflects the transferability between models (Goodfellow et al., 2014; Liu et al., 2016; Demontis et al., 2019), we early stop both federated and centralized models. For CIFAR10, we early stop both models at 90% of accuracy (75% for CNN). For ImageNet200, we follow Section 4.1 and early stop at 55%. We follow this training setting for the rest of this paper.
Tab. 2 shows that the adversarial examples generated by the federated model are highly transferable to both the federated and centralized model while adversarial examples generated by the centralized model exhibit less transferability. The T.Rate of federated-to-centralized attack is even larger than centralized-to-centralized attack. Secondly, T.Rate of adversarial examples between models trained under same paradigms is larger than models trained under different paradigms, which can be attributed to the difference of the two training paradigms, e.g., the discrepancy in the decision boundary (Goodfellow et al., 2014; Liu et al., 2016) or different sub-space (Tramèr et al., 2017).
@@ 4.3 Transfer Attack with limited data
In this section, we comprehensively evaluate the practicality and plausibility of our proposed attack setting, i.e. transfer attack with limited data. We present the overview of our attack setup in Algorithm 1. To simulate this scenario, we fix the generated partition used in the federated training and randomly select a specified number of users as malicious clients whose data is available for performing the attack. To perform the transfer attack, we train a surrogate model in a centralized manner with the collected data. Training details are specified in Section 3.1.
Input: A set of $N$ clients $\{C_{1},C_{2},\dots,C_{N}\}$ where $M$ clients $\{C_{1},C_{2},\dots,C_{M}\}$ are corrupted by attacker, datasets $X_{i}$ for each client $C_{i}$, sampling rate $K$, total communication round $T$, local iteration number $E$ and surrogate model training iteration number $T_{s}$.
Output: Adversarial perturbation $\delta$ on example $x$
/* Normal Federated Training with corrupted clients acting as benign */
initialize FL model $f_{0}$; for each round $t=1,2,\cdots,T$ do
       Server samples $K$ devices $S_{t}$ and distributes the global model $f_{t}$
      for each client $C_{i}$ in $S_{t}$ do in parallel
             $f_{t+1}^{i}\leftarrow\text{ClientUpdate}(f_{t},X_{i})$
       end forpar
      $f_{t+1}\leftarrow Aggregate(f_{t+1}^{1},\cdots,f_{t+1}^{K})$
end for
/* Train surrogate model */
Adversary collects data from corrupted clients $\{C_{1},C_{2},\dots,C_{M}\}$; initialize surrogate model $f^{\prime}_{0}$;
for each iter $t=1,2,\cdots,T_{s}$ do
       $f^{\prime}_{t}\leftarrow\text{ClientUpdate}(f^{\prime}_{t-1},\bigcup_{i=1,\cdots,M}X_{i})$
end for
$\delta\leftarrow\text{AdvPerturb}(f^{\prime},x)$;
Algorithm 1 Our Attack as Benign Setting
One of the key differences between the proposed setting and the conventional transfer-based attack is the amount of available data to train the substitute (source) model, which is a key factor for a successful attack since one would reason that more data will lead to a higher success rate. This is measured by the number of clients used to train the source model. To provide an overview of the transferability of adversarial examples generated by the source model trained with different numbers of clients, we plot their relation in Fig. 1. We have the following observations:
@@
From observation 2, we can see that the proposed attack can achieve comparable or even better T.Rate or lower T.Acc. Consequently, we can conclude that the proposed attack setting with limited data is likely to cause significant security breaches in the current and future FL applications. To further explain observation 1 and an intriguing phenomenon that the T.Rate of ResNet50 model rises to the peak and then decreases, we provide the following hypothesis: When the number of clients used to train the source model is small, the clean accuracy of the source model is also low, leading to a large discrepancy in the decision boundary. Increasing the number of users used in the source model minimizes such discrepancy until the amount of data is sufficient to train a source model with similar accuracy. At this point, the difference between the federated and centralized (Caldarola et al., 2022) becomes the dominant factor affecting the transferability since the source model is trained in the centralized paradigm.
Figure 1: Attack with data from a limited number of users. (a) We show the transfer rate of our attack with ResNet50 on the CIFAR10 dataset. We additionally experiment with attackers of different architectures. (b) experiment with CNN on CIFAR10 dataset. (c) we boost the performance of our attack with standard augmentation and pretraining techniques. (d) we perform our attack on different training stages. (e) we provide experiments on easier scenarios with 10 and 30 users, which further demonstrate the threat our attack poses.
@@
Training surrogate model in a federated manner. To validate the conjecture above, we train the surrogate model in a federated manner. Specifically, since we know each sample’s client, we partition the collected data from malicious clients as the federated model and train the source model in a federated manner. Results can be seen in (c) of Fig. 1: with a federated source model, the T.Rate can be slightly boosted at the beginning (limited number of clients) but continue to increase as the number of users increases and finally contributes to a significantly high T.Rate of 99%. This demonstrates our conjecture and also shows that if the hostile party trains the surrogate model in a federated manner, the T.Rate can be further increased. We further evaluate whether the partition information is crucial to the higher transferability in Appendix A.
@@
Boosting transfer rate. Following the conjectures above, we propose several effective approaches to enhance the transferability of the surrogate model trained by the malicious data only. We first boost the T.Rate of the source model trained with limited data with data augmentation and model pretraining. Without loss of generality, we leverage AutoAugment (Cubuk et al., 2018) and ImageNet pretrain. We believe other forms of augmentation and pretrained weights will exhibit similar effects. From (c) of Fig. 1, we can see, with these techniques we can successfully increase the T.Rate of 1% and 2% of clients from around 3% to 36% and 48% respectively. With 7% or 10% of clients’ data, the proposed attack setting achieves a high T.Rate of more than 80% (10% higher than the transfer-based attack). With simple training techniques, malicious clients can attack with more than 40% success rate with one or more clients and 80% with 7 to 10 clients.
Advanced Transfer Attack. To further show the threat posed by our attack, we leverage more advanced attacks to replace the default PGD attack as shown in (c) of Fig. 1: both AutoAttack (Croce & Hein, 2020) and skip attack (Wu et al., 2020a) achieve a transfer rate of over 80% with only 3 and 10 users respectively. Noticeably, AutoAttack achieves the highest transfer rate of 85% with 10% of the users.
Comparison with query-based attack As elaborated in Sec. 2, our transfer attack with limited data is similar but different from the query-based black-box attack, as we assume the malicious clients possess a portion of the original training data. Query-based black-box attack, on the other hand, aims to attack the target model with limited queries to the target model. Through these queries, the source model manages to optimize the decision boundary towards the target model. However, most APIs and FL-based systems require charges to access the service or equip anomaly detection that detects multiple or malicious queries.
We provide a comparison of our proposed attack setting and the query-based black-box attack. To facilitate a fair comparison, we follow the same experiment setting to train the query-based model. We plot the comparison in (a) of Fig. 2. Despite query-based outperforming the default of our proposed attack by a slight margin at the cost of expensive queries, our attack combined with federated training outperforms the query-based attack consistently at all configurations. It is also pertinent to note that facilitating the queries for data from only one client necessitates the execution of 500 requests to the target model (in a 100-client partition). Further, the query cost demonstrates a proportional escalation as the number of required queries increases, which may be regarded as impracticable in tangible operational contexts. Moreover, we underscore that our attack does not counteract the query-based attack. In fact, we can perform both if the FL system allows a certain number of queries. We believe that, combined with a limited number of queries, our attack strategy will be augmented to attain a higher transfer rate since the queries from the target model can help the surrogate model learn a decision boundary that more congruently aligns with that of the target model.
Figure 2: (a) Comparison with query-based black-box attack. (b) Attack with the surrogate model trained with the same distribution (no participation in the FL training) (c) Attack with the surrogate model trained with similar knowledge, i.e. CIFAR100.
@@
Practical evaluation. To practically evaluate our attack setting, we consider two cases: 1. the malicious clients can only participate in some periods during the FL training and 2. the attacker has no knowledge of the model deployed (i.e. unknown architecture). We simulate these situations by attacking in different training stages and with different architectures, as shown in (d) of Fig. 1.
Is Data Distribution or Similar Knowledge Sufficient? We emphasize the significance of the process that malicious clients stay benign during training and affect the training with their own data. We demonstrate in this section that this obtained training data is crucial to a successful attack.
We first show that using a similar distribution is not as good as the actual training data to train a surrogate model, as shown in (b) of Fig. 2. We note that the inherent heterogeneity of FL training will further impose challenges. The distinct distribution of each client’s data makes it nearly impossible to approximate an overall training distribution due to the scale and variability of the data.
Training the surrogate model with a dataset of similar knowledge or characteristics is also ineffective. We simulate this by using the CIFAR100 dataset to train a surrogate model and perform the attack against the FL model trained on the CIFAR10 dataset, as shown in (c) of Fig. 2. We also want to emphasize that our approach to obtaining the training data by acting benignly during training is practically impossible to defend against as there is no way to distinguish a benign client and a malicious client if they stay benign.
More Configurations. We also conduct experiments with 10 and 30 users as shown in (e) of Fig. 1 which is a relatively simpler setting. Specifically, we emphasize that when the federated model is trained on 10 and 30 uses, our attack achieves the highest 90% transfer rate with only 2 and 5 malicious clients respectively. When there are only 1 malicious client in the 10-user setting and 3 malicious clients in the 30-user setting, we can achieve over 60% transfer rate and over 70% transfer rate. We further perform experiments on CIFAR100, SVHN and a much larger and more realistic dataset ImageNet200, as demonstrated in (a) of Fig. 5 where similar trends are demonstrated. These results emphasize the threat of our attack setting.
@@ 5 Two intrinsic properties contributing to transfer robustness
To fully understand how adversarial examples transfer between centralized and federated models, we study two intrinsic properties of FL and its relation with transfer robustness. To protect the privacy of clients and leverage the massive data from user-end, FL utilizes distributed data to train a global model through local updates and aggregation at the server (McMahan et al., 2017). As a consequence, the heterogeneity of the distributed data and the aggregation operation is the core component of an FL method. In this section, we study how these two properties affect the transfer robustness of the FL model.
Figure 3: T.Rate vs. data of different heterogeneity and dispersion degree. (a): top 3 are results of ResNet50; (b): bottom 3 are results of CNN; Left: T.Rate as a function of the number of users in federated training; Middle: T.Rate as a function of dirichlet alpha; Right:T.Rate as a function of unbalanced sgm.
@@ 5.1 Decentralized training and data Heterogeneity
This section aims to explore the relationship between the degree of heterogeneity and transfer robustness.
@@
Control the degree of dispersion and heterogeneity. To explore the impact of distributed data on adversarial transferability, we control decentralization and heterogeneity through four indexes. By varying the number of clients in the partition, we alter the degree of dispersion of distributed data. We provide two approaches to control the heterogeneity of the distributed data. 1. Change the number of maximum classes per client (McMahan et al., 2017). 2. Change the alpha value of Dirichlet distributions ($\alpha$) (Wang et al., 2020; Yurochkin et al., 2019) (smaller $\alpha$ means a more non-iid partition) and the log-normal variance (sgm) of the Log-Normal distribution (larger variance denotes more unbalanced partitions) used in unbalanced partition (Zeng et al., 2021). We leverage the FedLab framework to generate the different partitions (Zeng et al., 2021). For simplicity, we leverage the centralized trained model as the source model for the rest of the experiments.
Degree of decentralization reduces transferability. We first explore the relation between decentralization and transferability. To control the degree of decentralization, we generate partitions with different numbers of clients and train target federated models on these partitions. Then we perform the transfer attack using the centralized model as the source model. As seen in Fig. 3 (left of (a) and (b)), we can observe that despite some fluctuations, T.Rate drops with an increasing number of users, which demonstrates that more decentralized data leads to lower transferability. We provide statistical testing for the correlation coefficient in Appendix C to further validate the above observation. With the Spearman correlation coefficient, we report a significant negative correlation on both ResNet50 and CNN between the degree of decentralization and T.Rate under a significance level of 0.1 with $p$-value=0.03 and 0.01 respectively for ResNet50 and CNN. We also visualize linear regression to fit the negative correlation between the degree of decentralization and T.Rate the in Appendix H.
Figure 4: Transfer rate v.s. maximum number of classes per client; (a): Results of ResNet50 on CIFAR10 dataset; (b): Results of CNN on CIFAR10 dataset; It is noteworthy to observe that the transfer rate in the 10-user setting persistently surpasses that in the 100-user setting. This further substantiates the proposition that more decentralized training leads to lower adversarial transferability for the federated model.
Data heterogeneity affects transfer attack. As discussed in Sec. 5.1, we provide two approaches for controlling the heterogeneity of the data. First, we alter the alpha values of the Dirichlet distributions to generate heterogeneous data of different degrees. As per the middle plot of (a) and (b) in Fig. 3, we can see that T.Rate increases as $\alpha$ increases.
We also alter the maximum number of classes per client to generate heterogeneous data of different degrees. Results are reported in Fig. 4. We can observe from Fig. 4 a clear increase trend in both the 10-user partition and the 100-user partition setting with ResNet50 and CNN. As the degree of heterogeneity decreases, the transferability increases. Experiments in both settings can illustrate our findings. This proves that our observations hold to wider circumstances and scenarios.
We further explore whether unbalanced data (i.e. different clients possess different numbers of samples) leads to a decrease in transferability in Fig. 3 (right of (a) and (b)). We can observe that a larger variance leads to a lower transfer rate, meaning that unbalanced data also contribute to higher robustness.
To validate the above observation, we provide statistical testing for the correlation coefficient in Appendix C. With the Spearman correlation coefficient, we report a significant negative correlation on ResNet50 between log-normal variance and T.Rate under a significance level of 0.1 ($p$-value=0.05). We report a significant correlation on all results with a level of 0.05 except the CNN experiments with different Dirichlet $\alpha$ and unbalanced sgm. Visualization of linear fitting is in Appendix H. To provide a more practical evaluation, we follow the setting in practical evaluation of Sec. 4.3 and evaluate the above findings in different training stages and with different architectures. As per (c) and (d) of Fig. 3, we can observe a similar correlation and trends between Dirichlet alpha and transfer rate. This further illustrates that our findings hold in various settings and configurations.